# Shaping Shared Languages: Human and Large Language Models’ Inductive Biases in Emergent Communication

### v1

@@ 4 Evaluation
We use linear mixed effect models to analyse our results. Specifically, we fit where can either be TopSim, RatioUniLabels and is the average value of two players in a round. To measure effects across conditions, we use . The slope determines the direction of the effect and the rate of change. Additionally, we use conditional , and marginal Nakagawa and Schielzeth (2013). The former considers fixed and random effects, to show how much variance can be explained by the model. Higher values of indicate that the model captures more variance and that correlations are stronger. , describes how much variance can be explained by the fixed effects. We report Pearson’s R to describe the relationship between TopSim and GenScore, and use a paired T-test, or Welch’s test when assumptions on normality and variance were not met, to assess whether metrics significantly differ.
@@ 5 Results
happened in a way that was expected based on earlier work Kirby et al. (2015); Raviv et al. (2021). The results of these 15 Human-Human () experiments act as a benchmark of human behaviour in our setup. We find that learning artificial languages is not a trivial task. After two rounds of exposure, labels were correctly guessed approximately half the time (). Freely labelling stimuli was done correctly only in of the labels. Nevertheless, reliable communication protocols emerged during the communicating block, interactions were significantly more successful in the final round compared to the first round (Figure 3, ). TopSim positively influenced communication (PercCom, ) and generalisation to new stimuli was more consistent when the languages that evolved during communication displayed more TopSim (). A qualitative inspection revealed that, during communication, participants quickly replaced the labels they learned before, with only parts of labels ’surviving’ this cut. Moreover, the number of shapes displayed (1, 2, or 3) was sometimes encoded by repeating the shape and colour labels several times, e.g., ”pufepufe” was used to indicate two green shapes. Although expressive, this solution does not generalise to larger numbers and is therefore arguably not compositional.
@@
was assessed by Kouwenhoven et al. (2025). Here, we briefly discuss the results of LLM-LLM () simulations.222We also ran simulations with other (smaller) models (i.e. Llama 3 8B, OLMo 2 7B 1124 & OLMo 2 13B 1124). Communication was successful but more difficult. OLMo 2 models struggled most. LLMs guessed labels correctly for , and the produced labels in the labelling block exactly matched the initial labels for . Indicating that learning the language is easier for LLMs than for humans. This is not surprising since the target stimulus is present in the prompt context in these blocks and there is virtually no memory constraint. Communication happens reliably as well (Figure 3), however, communication can–but does not always–result in degenerate vocabularies with few uniquely used labels (), significantly differing from human diversity in labels (, Figure 4). Interestingly, this happens even though the ratio of unique labels is modestly related to PercCom during communication (), suggesting that expressiveness is beneficial. A tentative explanation could be that aligning vocabularies happens much faster than in the other conditions. While humans would optimise languages whilst retaining expressiveness, LLMs start producing more duplicate and longer labels.
@@
Our main contribution comes from the Human-LLM condition in which participants () collaborated with LLMs. Successful communication necessitates that both entities adopt their behaviours and that a reliable referentially grounded vocabulary emerges. This is especially interesting since we observed that the abilities to learn artificial languages differ between humans and LLMs, and that the optimised languages differ in their use of homonyms, i.e., duplicate labels for different meanings. Despite these differences, communication is still possible (, Figure 3). The final performance was lower than the other conditions but the data suggests that prolonged interactions may result in higher communicative success. This exciting result shows that, even though the learning mechanisms of both entities may initially learn different signal-meaning relations, communication is possible. As such, the process of repeated learning and use of these artificial languages can overcome initial differences and indeed shape languages to be optimised for humans and LLMs. Out of 15 participants, 9 believed their partner was another human, despite communicating with an LLM. Performance did not change significantly as a result of this.
@@
Having established that participants can communicate, we now examine if and how languages differ across different experimental conditions, focussing on the language metrics. Generally, we find that languages optimised for LLMs differ from those optimised for humans, and that languages optimised for both are more human-like than LLM-like (Figure 4). This is most notably visible in the ratio of uniquely produced labels, their respective lengths, and the Ngram diversity. Languages optimised for humans contain more unique labels, have higher Ngram diversity, and are shorter when compared to languages optimised for LLMs. The latter of which corroborates the well-known human preference for efficient communication Smith and Culbertson (2020). The length, number of unique labels, and diversity of label parts resulting from human-LLM collaboration seem to adhere more to human preferences than to LLM preferences. A similar pattern is visible for the compositionality metrics that allow for variation. There is more homonymy in LLM-optimised languages than in human-LLM languages, suggesting that the meanings of these words should be disambiguated by the context (i.e., the distractor stimuli) in which they appeared. This strategy is not straightforward and perhaps requires more cognitive capacity than available for humans, and may therefore be lower in the collaborative condition. It moreover seems that humans introduce synonymy into the languages, i.e., they use more than a single character to refer to specific attribute values. This introduces variability that can explain why the ratio of unique labels and Ngram diversity is higher in the collaborative condition. Finally, the word order of messages is somewhat flexible for humans, while LLMs converge to a more fixed word order. The languages shaped by both seem to have human-like word order freedom. While these variations may introduce difficulties for LLMs to decode the meanings during communication, we do not find that PercCom is affected by the degree of homonymy, synonymy, or freedom.
@@ 6 Discussion
The primary goal of our work was to investigate if and how artificial languages differ when they are optimised for human and artificially intelligent language users. To do so, we extended work suggesting that LLMs can shape and use languages in referential communication Kouwenhoven et al. (2025). Their setup was adapted so that participants could interact with other human participants and with LLMs. This enabled controlled comparisons between the languages that evolve under different conditions. Our findings showed that human pairs, LLM pairs and Human-LLM pairs can learn and successfully use languages in a referential game. This suggests that mechanisms that influence how language evolves for humans, specifically, learning and using a language repeatedly (Smith, 2022, e.g.), also apply to computational and collaborative Human-LLM settings. In all conditions, successful communication was achieved by optimising an initially holistic unstructured vocabulary, to fit better with the inductive biases of the language users. We compared the languages across conditions and revealed that 1) while very human-like, LLM languages tend to be more strict (i.e., there is little variation) and that 2) languages adapted for human-LLM pairs tend to be more human-like than LLM-like (i.e., they are more diverse and have variation). Overall, our findings corroborate earlier claims that interactions between humans and machines are beneficial to establishing some form of referential grounding Mikolov et al. (2018); Kouwenhoven et al. (2022); Beuls and Van Eecke (2024).
@@
On the level of vocabulary, the ratio of uniquely produced labels by LLMs revealed that vocabularies can become degenerate. While this is also observed in human experiments when there is no pressure against it Kirby et al. (2008), communicative success as a pressure is typically enough to prevent this (Kirby et al., 2015, e.g.). In contrast, even though the LLMs in our experiments were tested in a communicative setting, this did not prevent the languages from becoming underspecified. Possibly this happened because the instruction to obtain communicative success was not explicit enough and did not induce enough pressure for expressivity. Or it may be that the distractors did not require the labels to be very specific but instead allowed using underlying concurrences that were picked up by LLMs but not by humans. This would also explain the wide range of scores on this RatioUniLabels. On a character level, we see related patterns in the form of high levels of homonymy, meaning that attribute values could be associated with multiple label characters, and that context was necessary to disentangle the correct meaning. While the duplicate labels can explain these scores for LLM simulations, this is not the case for humans. Here, the surprising behaviour of repeating label parts to indicate the amount attribute can explain the homonymy values.
@@
The process of tokenisation plays an important role in these simulations. One could argue that it may help in learning some mapping between tokens and meanings. The meaning attributes and their values are common English words while the initialised artificial languages consist of non-words that are tokenised into separate tokens. Meaning that the LLM is presented with a parsed set of attribute meanings and chunks of labels (i.e., the tokens). All that is left, is to attend to a specific token given a specific meaning, which is exactly what a transformer model is made for. Nevertheless, this does not undermine that these models indeed attribute attention correctly and that this produces human-like languages.
@@
Similarly, our observation that a shared referential communicative system can be established is quite remarkable. Humans and LLMs may well use completely different mechanisms, and learn different relations between meanings and signals. Yet, their vocabularies become referentially grounded and are pragmatically understood by both humans and LLMs. This confirms that even though LLMs are not trained for this task, they can be used as relatively unbiased language learners Wilcox et al. (2023), thereby providing a concrete example of how a pragmatic view of understanding as argued for by van Dijk et al. (2023) can be beneficial for collaborative tasks. This work also underscores the point made by Millière and Rathkopf (2024) that how LLMs or other AI models solve a cognitive task cannot be used as an argument against particular cognitive competences or language understanding, as long as the solution generalises.
@@
Our results concretely corroborate the idea that insights from emergent communication literature can inform and improve language learning in language models Smith et al. (2024); Beuls and Van Eecke (2024); Galke and Raviv (2025). We observed that just as languages accommodate for specific abilities and preferences in humans Josserand et al. (2024), Human-LLM languages also adapt to the abilities and preferences of their users in that they are more human-like than LLM-like. Specifically, human preferences for simplicity and efficiency Kirby et al. (2015); Gibson et al. (2019) likely drove vocabulary diversity while reducing lengths to human-like levels. This indicates that, in this experiment, LLMs are more flexible communicative partners than humans. These findings reinforce the idea that repeated interactions with humans are crucial to maintaining referentially grounded human-like vocabularies instead of training only on recursively generated data (Shumailov et al., 2024, e.g.) or using AI-augmented optimisation algorithms (Lee et al., 2024, e.g.). In particular, this underscores the need for new methods that include human interaction in the training processes of LLMs and shows that using communicative success as a reward signal can be a fruitful, novel direction.
@@
Finally, we acknowledge that our results depend on methodological considerations such as the use of in-context learning, the prompt format, and the sampling method. However, the primary goal was to extend previous work by investigating if languages optimised for human and LLM preferences can evolve. As such we stayed close to well-known experimental methods in the field of language emergence and used prompts developed in Kouwenhoven et al. (2025). Importantly, we did not optimise for communicative success, human-like results, or compositional vocabularies.

### v2

@@ 4 Evaluation
We use linear mixed effect models to analyse our results. Specifically, we fit where can either be TopSim, RatioUniLabels and is the average value of two players in a round. To measure effects across conditions, we use . The slope determines the direction of the effect and the rate of change. Additionally, we use conditional , and marginal Nakagawa and Schielzeth (2013). The former considers fixed and random effects to show how much variance can be explained by the model. Higher values of indicate that the model captures more variance and that correlations are stronger. describes how much variance can be explained by the fixed effects. We report Pearson’s R to describe the relationship between TopSim and GenScore, and use a paired T-test, or Welch’s test when assumptions on normality and variance were not met, to assess whether metrics significantly differ.
@@ 5 Results
happened in a way that was expected based on earlier work Kirby et al. (2015); Raviv et al. (2021). The results of these 15 Human-Human () experiments act as a benchmark of human behaviour in our setup. We find that learning artificial languages is not a trivial task. After two rounds of exposure, labels were correctly guessed approximately half the time (). Freely labelling stimuli was done correctly only in of the labels. Nevertheless, reliable communication protocols emerged during the communicating block, interactions were significantly more successful in the final round compared to the first round (Figure 3, ). TopSim positively influenced communication (PercCom, ) and generalisation to new stimuli was more consistent when the languages that evolved during communication displayed more TopSim (). A qualitative inspection revealed that, during communication, participants quickly replaced the labels they learned before, with only parts of labels ‘surviving’ this cut. Moreover, the number of shapes displayed (1, 2, or 3) was sometimes encoded by repeating the shape and colour labels several times, e.g., ”pufepufe” was used to indicate two green shapes. Although expressive, this solution does not generalise to larger numbers and is therefore arguably not compositional.
@@
was assessed by Kouwenhoven et al. (2025). Here, we briefly discuss the results of LLM-LLM () simulations 222We also ran simulations with other (smaller) models (i.e. Llama 3 8B, OLMo 2 7B 1124 & OLMo 2 13B 1124). Communication was successful but more difficult. OLMo 2 models struggled most.. LLMs correctly guessed labels for , and the produced labels in the labelling block exactly matched the initial labels for . This indicates that learning the language is easier for LLMs than for humans. This is not surprising since the target stimulus is present in the prompt context in these blocks and there is virtually no memory constraint. Communication happens reliably as well (Figure 3), however, communication can–but does not always–result in degenerate vocabularies with few uniquely used labels (), significantly differing from human diversity in labels (, Figure 4). Interestingly, this happens even though the ratio of unique labels is modestly related to PercCom during communication (), suggesting that expressiveness is beneficial. A tentative explanation could be that aligning vocabularies happens much faster than in the other conditions. While humans would optimise languages whilst retaining expressiveness, LLMs start producing more duplicate and longer labels.
@@
Our main contribution comes from the Human-LLM condition in which participants () collaborated with LLMs. Successful communication necessitates that both entities adopt their behaviours and that a reliable referentially grounded vocabulary emerges. This is especially interesting since we observed that the abilities to learn artificial languages differ between humans and LLMs, and that the optimised languages differ in their use of homonyms, i.e., duplicate labels for different meanings. Despite these differences, communication is still possible (, Figure 3). The final performance was lower than the other conditions, but the data suggests that prolonged interactions may result in higher communicative success. This exciting result shows that, even though the learning mechanisms of both entities may initially learn different signal-meaning relations, communication is possible. As such, the process of repeated learning and use of these artificial languages can overcome initial differences and indeed shape languages to be optimised for humans and LLMs. Out of 15 participants, 9 believed their partner was another human, despite communicating with an LLM. Performance did not change significantly as a result of this.
@@
Having established that participants can communicate, we now examine if and how languages differ across different experimental conditions, focusing on the language metrics. Generally, we find that languages optimised for LLMs differ from those optimised for humans, and that languages optimised for both are more human-like than LLM-like (Figure 4). This is most notably visible in the ratio of uniquely produced labels, their respective lengths, and the Ngram diversity. Languages optimised for humans contain more unique labels, have higher Ngram diversity, and are shorter when compared to languages optimised for LLMs. The latter of which corroborates the well-known human preference for efficient communication Smith and Culbertson (2020). The length, number of unique labels, and diversity of label parts resulting from human-LLM collaboration seem to adhere more to human preferences than to LLM preferences. A similar pattern is visible for the compositionality metrics that allow for variation. There is more homonymy in LLM-optimised languages than in human-LLM languages, suggesting that the meanings of these words should be disambiguated by the context (i.e., the distractor stimuli) in which they appeared. This strategy is not straightforward and perhaps requires more cognitive capacity than available for humans, and may therefore be lower in the collaborative condition. It moreover seems that humans introduce synonymy into the languages, i.e., they use more than a single character to refer to specific attribute values. This introduces variability that can explain why the ratio of unique labels and Ngram diversity is higher in the collaborative condition. Finally, the word order of messages is somewhat flexible for humans, while LLMs converge to a more fixed word order. The languages shaped by both seem to have human-like word order freedom. While these variations may introduce difficulties for LLMs to decode the meanings during communication, we do not find that PercCom is affected by the degree of homonymy, synonymy, or freedom.
@@ 6 Discussion
The primary goal of our work was to investigate if and how artificial languages differ when optimised for human and artificially intelligent language users. We extended work suggesting that LLMs can shape and use languages in referential communication Kouwenhoven et al. (2025). Their setup was adapted so that participants could interact with other human participants and with LLMs. This enabled controlled comparisons between the languages that evolve under different conditions. Our findings showed that human pairs, LLM pairs and Human-LLM pairs can learn and successfully use languages in a referential game. This suggests that mechanisms that influence how language evolves for humans, specifically, learning and using a language repeatedly (Smith, 2022, e.g.), also apply to computational and collaborative Human-LLM settings. In all conditions, successful communication was achieved by optimising an initially holistic unstructured vocabulary, to fit better with the inductive biases of the language users. Comparison of the languages across conditions revealed that 1) while very human-like, LLM languages tend to be more strict (i.e., there is little variation) and that 2) languages adapted for human-LLM pairs tend to be more human-like than LLM-like (i.e., they are more diverse and have variation). Overall, our findings corroborate earlier claims that interactions between humans and machines are beneficial to establishing some form of referential grounding Mikolov et al. (2018); Kouwenhoven et al. (2022); Beuls and Van Eecke (2024).
@@
On the level of vocabulary, the ratio of uniquely produced labels by LLMs revealed that vocabularies can become degenerate. While this is also observed in human experiments when there is no pressure against it Kirby et al. (2008), communicative success as a pressure is typically enough to prevent this (Kirby et al., 2015, e.g.). In contrast, even though the LLMs in our experiments were tested in a communicative setting, this did not prevent the languages from becoming underspecified. Possibly because the instruction to obtain communicative success was not explicit enough and did not induce enough pressure for expressivity. Or it may be that the distractors did not require the labels to be very specific but instead allowed using underlying concurrences that were picked up by LLMs but not by humans. This would also explain the wide range of scores on this RatioUniLabels. On a character level, we see related patterns in the form of high levels of homonymy, meaning that attribute values could be associated with multiple label characters, and that context was necessary to disentangle the correct meaning. While the duplicate labels can explain these scores for LLM simulations, this is not the case for humans. Here, the surprising behaviour of repeating label parts to indicate the amount attribute can explain the homonymy values.
@@
Since we do not modify the tokeniser to deal with our artificial languages, the tokenisation process plays an important role in these simulations. One could argue that this helps in learning a mapping between tokens and meanings. The meaning attributes and their values are common English words, while the initialised artificial languages are non-words tokenised into separate tokens. Meaning that the LLM is presented with a parsed set of attribute meanings and chunks of labels (i.e., the tokens). All that is left is to attend to a specific token given a specific meaning attribute, which is exactly what a transformer model is made for. This does, however, not undermine that these models indeed attribute attention correctly and that this produces human-like languages.
@@
The fact that a shared referential communicative system can be established is quite remarkable. Humans and LLMs may well rely on completely different mechanisms, and learn different relations between meanings and signals. Yet, their vocabularies become referentially grounded and are pragmatically understood by both humans and LLMs. This confirms that even though LLMs are not trained for this task, they can be used as relatively unbiased language learners Wilcox et al. (2023), providing a concrete example of how a pragmatic view of understanding as argued for by van Dijk et al. (2023), can be beneficial for collaborative tasks. This work also underscores the point made by Millière and Rathkopf (2024) that how LLMs or other AI models solve a cognitive task cannot be used as an argument against particular cognitive competences or language understanding, as long as the solution generalises.
@@
Our results concretely corroborate the idea that insights from emergent communication literature can inform and improve language learning in language models Smith et al. (2024); Beuls and Van Eecke (2024); Galke and Raviv (2025). We observed that just as languages accommodate for specific abilities and preferences in humans Josserand et al. (2024), Human-LLM languages also adapt to the abilities and preferences of their users in that they are more human-like than LLM-like. Specifically, human preferences for simplicity and efficiency Kirby et al. (2015); Gibson et al. (2019) likely drove vocabulary diversity while reducing lengths to human-like levels. This indicates that, in this experiment, LLMs are more flexible communicative partners than humans. These findings reinforce the idea that repeated interactions with humans are crucial to maintaining referentially grounded human-like vocabularies instead of training only on recursively generated data (Shumailov et al., 2024, e.g.) or using AI-augmented optimisation algorithms (Lee et al., 2024, e.g.). In particular, this underscores the need for new LLM training methods that include human interaction and shows that using communicative success as a reward signal can be a fruitful, novel direction (Stöpler et al., 2025, e.g.)
@@
Finally, we acknowledge that our results depend on methodological considerations such as in-context learning, the prompt format, and the sampling method. However, the primary goal was to extend previous work by investigating if languages optimised for human and LLM preferences can evolve. As such, we stayed close to well-known experimental methods in language evolution research and used prompts developed in Kouwenhoven et al. (2025). Importantly, we did not optimise for communicative success, human-like results, or compositional vocabularies.